OpenStudy (anonymous):

Write the equation of the quadratic function with roots -10 and -8 and a vertex at (-9, -3).

hero (hero):

Satellite pretty much showed you the format for this

OpenStudy (anonymous):

is it the same for all?

hero (hero):

Pretty much

OpenStudy (anonymous):

still stumped...O.o

OpenStudy (anonymous):

but ok

OpenStudy (anonymous):

Here's what you can do with this information. If the quadratic has a root of x = k, then it has a factor of (x-k). i.e. for roots x = -10 and x = -8, the quadratic has factors (x+10)(x+8). There may still be an additional constant factor, a, so simplify y = a(x+10)(x+8) to get it in the form, ax^2+bx+c. With the additional information of a vertex at (-9, -3), then you can use the vertex form of the equation y=a(x-h)+k and put in your given information. The vertex is at (h, k), so y = a(x+9)^2 - 3. If you simplify that, you get y = ax^2 +18ax +81a -3. Compare that with the other equation to solve for a.